# Spirals and Beyond: Competitive Plane Search with Multi-Speed Agents

### v1

@@ 4.1 The Speed-Dependent Off-Set Spiral Algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$
In this section, we generalize the trajectory $\{f_{i}^{k}\}_{i\in[n]}$ for $\textsc{MSP}_{n}(\textbf{1})$ in order to propose a polynomial-time algorithm that solves $\textsc{MSP}_{n}(c)$ for arbitrary $c\in\mathbb{R}^{n}$. Recall that we made the assumptions about $\{c_{i}\}_{i\in[n]}$ are given in non-decreasing order and that $c_{0}=1$. As a reminder, we also refer to agent-0 as agent-$n$, and we use $c_{0}=c_{n}=1$. In that direction, we propose a family of spiral-type search algorithms, as described in Algorithm 2, which we call Off-Set Trajectories, in which agents' trajectories do not intersect. The key difference from Algorithm 1 is that agents in the new trajectories are assigned arbitrary angular separations, parameterized by values $\phi_{i}$, which will later be defined in terms of the agents' speeds. These degrees of freedom, in addition to the spiral expansion parameter $k$, enable a competitive analysis that is agent-independent, a defining optimality property, thereby improving the overall efficiency of the agents that will contribute to search.
@@
Note that the parameters $\phi_{0}=2\pi$ and $\phi_{n}=0$ are fixed and correspond to agent-0 (also referred to as agent-$n$ for notational convenience). Due to the periodicity of trigonometric functions, we have $g_{0}(t)=g_{n}(t)$, ensuring cyclic behavior in the agents' trajectories. As before, the condition $k>0$ is necessary for the correctness of the search algorithm, guaranteeing that every point in the plane is eventually exposed. The ordering $\phi_{0}\geq\phi_{1}\geq\dots\geq\phi_{n-1}\geq 0$ is a normalization that determines the appropriate sequence in which agents should be considered based on their speeds. This ordering ensures that whenever agent $i$ reaches a given ray, the next agent to do so is agent $(i+1)\bmod n$. Consequently, every ray is visited in cyclic order by agents $0,1,2,\dots,n-1$, repeating indefinitely.
@@
To compute these parameters, we propose Algorithm 3, which utilizes precomputed expansion factors $\kappa_{j}$ from Lemma 3.4 and the search cost $\mathcal{U}_{j}$ for $j$ unit-speed agents from Theorem 2.1. Algorithm 3 returns parameters $\phi_{j}$ only for the agents that actively contribute to the search. Specifically, it identifies a key agent and all agents at least as fast, while ignoring slower agents, as they do not expose new points. The values of $\phi_{j}$ for these ignored agents are irrelevant to the solution.
@@ 4.2 Properties of the Off-Set Trajectories
In this section, we identify and analyze key properties of Algorithm 2. Notably, in this section we do not assume that the agent speeds $c_{i}$ are given in non-decreasing order. We begin by showing that agents following the trajectories $\{g_{i}^{k}(t)\}_{i\in[n]}$ search the plane in an invariant pattern.
Lemma 4.2.
@@
Proof.
Fix some $t_{0}\in\mathbb{R}$, and agent $i$ with speed $c_{i}$. We observe that the angular distance between points $g_{0}^{k}(t_{0})$ and $g_{i}^{k}(t_{0})$ equals $2\pi-\phi_{i}$. Hence, the two agents preserve their angular distance if and only if they reach the corresponding points $g_{0}^{k}(t_{0})$ and $g_{i}^{k}(t_{0})$ at the same time. We argue next that the length of trajectory $g^{k}_{i}$ up to $t=t_{0}$ is $\ell_{i}(t_{0})=c_{i}\ell_{0}(t_{0})$. Indeed, let $h:\mathbb{R}\to\mathbb{R}^{2}$ be as in the statement of Lemma 3.2. We note that $h(t)=g_{0}^{k}(t)$, since $\phi_{0}=2\pi$ and $c_{0}=1$. But then, by the same lemma, and for any speed-$c_{i}$, we have
@@
Proof.
Consider $i\in[n]$ and agents $i$ and $i+1$ with speeds $c_{i}$ and $c_{i+1}$, respectively, so that $\phi_{i}\geq\phi_{i+1}$. Let $t_{0}$ be the time at which agent-$i$ reaches $\mathcal{L}_{0}$. By rotating the plane around the origin, we may assume that $t_{0}+\phi_{i}=2\pi s$ for some $s\in\mathbb{Z}$. Using Lemma 4.2, we see that by time $\ell_{0}(t_{0})$, all points in the closed interval $0,c_{i}e^{kt_{0}}]=[0,c_{i}e^{2\pi sk-k\phi_{i}}]$ are already exposed by agent-$i$. Since time increases with $t_{0}$ and $\phi_{i}\geq\phi_{i+1}$, it follows that $0\leq\phi_{i}-\phi_{i+1}\leq 2\pi.$ The next time agent-$(i+1)$ reaches the positive horizontal axis is when it arrives at
@@
New points are exposed precisely when $c_{i+1}e^{2\pi ks-k\phi_{i+1}}>c_{i}e^{2\pi ks-k\phi_{i}},$ or equivalently, $c_{i+1}e^{-k\phi_{i+1}}>c_{i}e^{-k\phi_{i}}.$ Furthermore, the newly exposed points lie in the semi-open (and semi-closed) interval as stated in the lemma. ∎
@@
Now, fix $i\in[n]$. By Lemma 4.3, agent-$(i+1)$ exposes all points $x\in\left(c_{i}e^{2\pi ks-k\phi_{i}},c_{i}e^{2\pi ks-k\phi_{i+1}}\right],$ where $c_{0}=1$, $\phi_{0}=2\pi$, and agent-$n$ is identified with agent-0. For notational convenience, denote this interval as $\mathcal{I}^{s}_{i}$. By Lemma 4.2, agent-$(i+1)$ exposes $(x,0)$ at time
@@
for all $j\in[n]$, which, by Lemma 4.3, is equivalent to all agents exposing new points. Expanding this inequality using (2), we obtain
@@
Let $S_{j}$ denote the agent-$j$ search cost and $S$ the Off-Set trajectory search cost. We show that $S=S_{j}$ if and only if the trajectory parameters satisfy (2). Since we have already established the validity of $\{\phi_{i}\}_{i\in[n]}$ and ensured that all agents expose new points, we apply Lemma 4.4 to obtain
@@
Lemma 4.4 also provides the formula for $S=S(k,n,c)$. Solving this equation for $\phi_{j}$ yields (2), as required. ∎
@@
Next, we verify that the parameters returned by Algorithm 3 satisfy the conditions required for the Off-Set Algorithm 2. Specifically, let $\mu,\nu,\{\phi_{i}\}_{i\in[\mu]}$ be the returned parameters, where the expansion factor used in Theorem 4.1 is $k=\kappa_{\nu}$. By Step 2, we have $\mu<l$, and by Step 1 and Lemma 4.5, the parameters $\kappa_{\nu}$ and $\{\phi_{i}\}_{i\in[\mu]}$ are valid for the Off-Set Algorithm 2 using a subset of $\nu$ agents, indexed by $\mu]$. By the same lemma, all agents $i\in[\mu]$ expose new points, and the agent-$i$ search cost is independent of $i$. Thus, applying Lemma 4.4 and using the definition of $\mathcal{U}_{\nu}$ from Theorem 2.1, we conclude that the induced search cost of the Off-Set algorithm matches the bound given in Theorem 4.1, as required. ∎
@@
A crucial aspect of Algorithm 3 is that it filters out agents with insufficiently fast speeds. Among all agents with ``sufficiently large'' speeds, it selects a subset of the strictly fastest agents, denoted as $A$, to minimize the ratio $\mathcal{U}_{|A|}/\mathcal{G}_{|A|}.$
An interesting observation is that Theorem 4.1 generalizes Theorem 2.1, whose proof also serves as a key building block in establishing the former result. In particular, consider $n$ unit-speed agents, meaning $c_{i}=1$ for all $i\in[n]$ is given as input to Algorithm 3. These agents are already sorted in non-decreasing order, amd their geometric means satisfy $\mathcal{G}_{i}=1$ for every $i\in[n]$. Thus, the largest index $l$ satisfying Step 1 of the algorithm is l=n1l=n-1, as $\mathcal{G}_{n}e^{-2\pi\kappa_{n}/n}=e^{-2\pi\kappa_{n}/n}<1,$ where the inequality holds because $\kappa_{n}>0$. By Proposition 3.5, the values $\mathcal{U}_{i}$ are strictly decreasing in $i$, so the smallest search cost, as determined by Step 2, is achieved at $\mu=n-1$. Then, in Step 3, the algorithm sets $\nu=\mu+1=n$, and the angular parameters are assigned as $\phi_{i}=\tfrac{2\pi(n-j)}{n},$ for $j\in[n]$, which exactly matches the definition in Algorithm 1.
@@ 4.4 A Simplified Algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$
This section is devoted to proving Theorem 2.2, that is, we describe a simpler algorithm than $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$ with the same cost search performance. For this, given input speeds $c\in\mathbb{R}^{n}$ with $c_{0}=1$ (recall we identify agent-0 as agent-$n$, and hence $c_{0}=c_{n}=1$) where the speeds are in non-decreasing order, we set
@@
Lemma 4.6.
Let $\kappa_{n}$ be as in Lemma 3.4. Set
@@
Regarding (a), we recall Lemma 3.5 according to which $\kappa_{n}$ is positive, strictly increasing, and strictly concave in $n$. Then, we observe that $A(x)$ is decreasing in $x>0$, as well as that $\lim_{x\rightarrow\infty}A(x)=1$. Moreover, by Lemma 3.4 it is easy to see that
@@ 5 Cone Search with a Unit-Speed Agent
In this section, we design a trajectory for searching for a point within an angle-$\phi$ cone, where $\phi\in(0,\pi)$, using a unit-speed agent. That is, we propose a solution to the problem $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$ and subsequently prove Theorem 2.4.
Figure 3: A depiction of the Bouncing Trajectory for $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$. Angle $\theta$ determines the sequence of points $\Gamma_{i}$, as per Lemma 5.2, using sequence $q_{i}=\left(\tfrac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}\right)^{i}$.
@@
Next, we compute the points $\Gamma_{i}$ in Algorithm 4 and derive the search cost using Lemma 5.1. The parameter $\beta=\beta(\phi,\theta)$ defines the sequence $\{g_{j}\}_{j}$, which uniquely determines the points $\{\Gamma_{j}\}_{j}$ and vice versa. We now provide a geometric and inductive definition of these points. For reference, Figure 3 illustrates the construction and highlights the role of the angle $\theta$.
@@
Using Lemma 5.1, the search cost of Algorithm 4 is
@@
By Lemma 5.2, if $\theta\in(\phi,\pi/2+\phi/2)$, the sequence is valid for the algorithm, and the resulting search cost is

### v2

@@ 4 Searching the Plane with $n$ Agents of Arbitrary Speeds
Our construction should be viewed as a best-possible extension of spiral-based search to heterogeneous speeds, under the strong requirement that all participating agents incur the same competitive cost. One of the goals of the paper is to understand whether this natural generalization remains optimal once agent speeds differ.
@@ 4.1 The Speed-Dependent Off-Set Spiral Algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$
In this section, we generalize the trajectory $\{f_{i}^{k}\}_{i\in[n]}$ for $\textsc{MSP}_{n}(\textbf{1})$ in order to propose a polynomial-time algorithm that solves $\textsc{MSP}_{n}(c)$ for arbitrary $c\in\mathbb{R}^{n}$. Recall that we made the assumptions about $\{c_{i}\}_{i\in[n]}$ are given in non-increasing order and that $c_{0}=1$. As a reminder, we also refer to agent-0 as agent-$n$, and we use $c_{0}=c_{n}=1$. In that direction, we propose a family of spiral-type search algorithms, as described in Algorithm 2, which we call Off-Set Trajectories, in which agents’ trajectories do not intersect. The key difference from Algorithm 1 is that agents in the new trajectories are assigned arbitrary angular separations, parameterized by values $\phi_{i}$, which will later be defined in terms of the agents’ speeds. These degrees of freedom, in addition to the spiral expansion parameter $k$, enable a competitive analysis that is agent-independent, a defining optimality property, thereby improving the overall efficiency of the agents that will contribute to search.
@@
For any fixed parameter value $t$, the distance of agent $i$ from the origin equals $c_{i}e^{kt}$, and thus the distances of any two agents remain proportional. Consequently, as $t$ varies, the point configuration $\{g_{i}^{k}(t)\}_{i\in[n]}$ evolves by uniform scaling and rotation, and any two such configurations are similar. This contrasts with the unit speed case, where all agents have the same distance from the origin at each snapshot, and the configuration forms a regular $n$-gon. Thus, in the multispeed setting, the configuration need not be regular, but its shape is preserved over time up to similarity, see Figure 3.
Figure 3: The depiction of trajectories of Algorithm 2. In this example, we use $n=3$ agents with speeds $c_{0}=1$, $c_{1}=0.9$, and $c_{2}=0.7$, and we plot the corresponding offset logarithmic spirals (as defined by the algorithm). We also show two snapshots of the agents, at two distinct parameter values, together with the triangles formed by connecting the three agent positions in each snapshot. In contrast to the unit speed case (Figure 2), where all agents lie on the same circle at every snapshot and hence form a regular 33-gon, the multispeed agents lie at different distances from the origin; these distances remain proportional, and the two triangles are similar (their side lengths scale by the same factor between snapshots).
Note that the parameters $\phi_{0}=2\pi$ and $\phi_{n}=0$ are fixed and correspond to agent-0 (also referred to as agent-$n$ for notational convenience). Due to the periodicity of trigonometric functions, we have $g_{0}(t)=g_{n}(t)$, ensuring cyclic behavior in the agents’ trajectories. As before, the condition $k>0$ is necessary for the correctness of the search algorithm, guaranteeing that every point in the plane is eventually exposed. The ordering $\phi_{0}\geq\phi_{1}\geq\dots\geq\phi_{n-1}\geq 0$ is a normalization that determines the appropriate sequence in which agents should be considered based on their speeds. This ordering ensures that whenever agent $i$ reaches a given ray, the next agent to do so is agent $(i+1)\bmod n$. Consequently, every ray is visited in cyclic order by agents $0,1,2,\dots,n-1$, repeating indefinitely.
@@
To compute these parameters, we propose Algorithm 3, which utilizes precomputed expansion factors $\kappa_{j}$ from Lemma 4 and the search cost $\mathcal{U}_{j}$ for $j$ unit-speed agents from Theorem 2.1. Algorithm 3 returns parameters $\phi_{j}$ only for the agents that actively contribute to the search. Specifically, it identifies a key agent and all agents at least as fast, while ignoring slower agents, as they do not expose new points. The values of $\phi_{j}$ for these ignored agents are irrelevant to the solution.
@@ 4.2 Properties of the Off-Set Trajectories
In this section, we identify and analyze key properties of Algorithm 2. Notably, in this section we do not assume that the agent speeds $c_{i}$ are given in non-increasing order. We begin by showing that agents following the trajectories $\{g_{i}^{k}(t)\}_{i\in[n]}$ search the plane in an invariant pattern.
Lemma 5
@@
Proof
Fix some $t_{0}\in\mathbb{R}$, and agent $i$ with speed $c_{i}$. We observe that the angular distance between points $g_{0}^{k}(t_{0})$ and $g_{i}^{k}(t_{0})$ equals $2\pi-\phi_{i}$. Hence, the two agents preserve their angular distance if and only if they reach the corresponding points $g_{0}^{k}(t_{0})$ and $g_{i}^{k}(t_{0})$ at the same time. We argue next that the length of trajectory $g^{k}_{i}$ up to $t=t_{0}$ is $\ell_{i}(t_{0})=c_{i}\ell_{0}(t_{0})$. Indeed, let $h:\mathbb{R}\to\mathbb{R}^{2}$ be as in the statement of Lemma 2. We note that $h(t)=g_{0}^{k}(t)$, since $\phi_{0}=2\pi$ and $c_{0}=1$. But then, by the same lemma, and for any speed-$c_{i}$, we have
@@
Proof
Consider $i\in[n]$ and agents $i$ and $i+1$ with speeds $c_{i}$ and $c_{i+1}$, respectively, so that $\phi_{i}\geq\phi_{i+1}$. Let $t_{0}$ be the time at which agent-$i$ reaches $\mathcal{L}_{0}$. By rotating the plane around the origin, we may assume that $t_{0}+\phi_{i}=2\pi s$ for some $s\in\mathbb{Z}$. Using Lemma 5, we see that by time $\ell_{0}(t_{0})$, all points in the closed interval $0,c_{i}e^{kt_{0}}]=[0,c_{i}e^{2\pi sk-k\phi_{i}}]$ are already exposed by agent-$i$. Since time increases with $t_{0}$ and $\phi_{i}\geq\phi_{i+1}$, it follows that $0\leq\phi_{i}-\phi_{i+1}\leq 2\pi.$ The next time agent-$(i+1)$ reaches the positive horizontal axis is when it arrives at
@@
New points are exposed precisely when
$c_{i+1}e^{2\pi ks-k\phi_{i+1}}>c_{i}e^{2\pi ks-k\phi_{i}},$
or equivalently,
$c_{i+1}e^{-k\phi_{i+1}}>c_{i}e^{-k\phi_{i}}.$
Furthermore, the newly exposed points lie in the semi-open (and semi-closed) interval as stated in the lemma. ∎
@@
Now, fix $i\in[n]$. By Lemma 6, agent-$(i+1)$ exposes all points $x\in\left(c_{i}e^{2\pi ks-k\phi_{i}},c_{i}e^{2\pi ks-k\phi_{i+1}}\right],$ where $c_{0}=1$, $\phi_{0}=2\pi$, and agent-$n$ is identified with agent-0. For notational convenience, denote this interval as $\mathcal{I}^{s}_{i}$. By Lemma 5, agent-$(i+1)$ exposes $(x,0)$ at time
@@
for all $j\in[n]$, which, by Lemma 6, is equivalent to all agents exposing new points. Expanding this inequality using (2), we obtain
@@
Let $S_{j}$ denote the agent-$j$ search cost and $S$ the Off-Set trajectory search cost. We show that $S=S_{j}$ if and only if the trajectory parameters satisfy (2). Since we have already established the validity of $\{\phi_{i}\}_{i\in[n]}$ and ensured that all agents expose new points, we apply Lemma 7 to obtain
@@
Lemma 7 also provides the formula for $S=S(k,n,c)$. Solving this equation for $\phi_{j}$ yields (2), as required. ∎
@@
Next, we verify that the parameters returned by Algorithm 3 satisfy the conditions required for the Off-Set Algorithm 2. Specifically, let $\mu,\nu,\{\phi_{i}\}_{i\in[\mu]}$ be the returned parameters, where the expansion factor used in Theorem 4.1 is $k=\kappa_{\nu}$. By Step 2, we have $\mu<l$, and by Step 1 and Lemma 8, the parameters $\kappa_{\nu}$ and $\{\phi_{i}\}_{i\in[\mu]}$ are valid for the Off-Set Algorithm 2 using a subset of $\nu$ agents, indexed by $\mu]$. By the same lemma, all agents $i\in[\mu]$ expose new points, and the agent-$i$ search cost is independent of $i$. Thus, applying Lemma 7 and using the definition of $\mathcal{U}_{\nu}$ from Theorem 2.1, we conclude that the induced search cost of the Off-Set algorithm matches the bound given in Theorem 4.1, as required. ∎
@@
A crucial aspect of Algorithm 3 is that it filters out agents with insufficiently fast speeds. Among all agents with “sufficiently large” speeds, it selects a subset of the strictly fastest agents, denoted as $A$, to minimize the ratio $\mathcal{U}_{|A|}/\mathcal{G}_{|A|}.$
An interesting observation is that Theorem 4.1 generalizes Theorem 2.1, whose proof also serves as a key building block in establishing the former result. In particular, consider $n$ unit-speed agents, meaning $c_{i}=1$ for all $i\in[n]$ is given as input to Algorithm 3. These agents are already sorted in non-increasing order, and their geometric means satisfy $\mathcal{G}_{i}=1$ for every $i\in[n]$. Thus, the largest index $l$ satisfying Step 1 of the algorithm is l=n1l=n-1, as $\mathcal{G}_{n}e^{-2\pi\kappa_{n}/n}=e^{-2\pi\kappa_{n}/n}<1,$ where the inequality holds because $\kappa_{n}>0$. By Proposition 1, the values $\mathcal{U}_{i}$ are strictly decreasing in $i$, so the smallest search cost, as determined by Step 2, is achieved at $\mu=n-1$. Then, in Step 3, the algorithm sets $\nu=\mu+1=n$, and the angular parameters are assigned as $\phi_{i}=\tfrac{2\pi(n-j)}{n},$ for $j\in[n]$, which exactly matches the definition in Algorithm 1.
@@ 4.4 A Simplified Algorithm for $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$
This section is devoted to proving Theorem 2.2, that is, we describe a simpler algorithm than $\mathcal{P}\text{-}\textsc{MSP}_{n}(c)$ with the same cost search performance. For this, given input speeds $c\in\mathbb{R}^{n}$ with $c_{0}=1$ (recall we identify agent-0 as agent-$n$, and hence $c_{0}=c_{n}=1$) where the speeds are in non-increasing order, we set
@@
Lemma 9
Let $\kappa_{n}$ be as in Lemma 4. Set
@@
Regarding (a), we recall Lemma 1 according to which $\kappa_{n}$ is positive, strictly increasing, and strictly concave in $n$. Then, we observe that $A(x)$ is decreasing in $x>0$, as well as that $\lim_{x\rightarrow\infty}A(x)=1$. Moreover, by Lemma 4 it is easy to see that
@@ 5 Cone Search with a Unit-Speed Agent
In this section, we design a trajectory for searching for a point within an angle-$\phi$ cone, where $\phi\in(0,\pi)$, using a unit-speed agent. That is, we propose a solution to the problem $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$ and subsequently prove Theorem 2.3. These results are developed not only for their independent geometric interest, but also to probe the limitations of spiral-based search when agents of different speeds are present. They will later be combined with the spiral-based constructions introduced earlier to obtain improved strategies for the multi-speed point-search problem, providing evidence that spiral-based trajectories, which are optimal for single-agent search, can be suboptimal in the multi-agent, multi-speed setting.
Figure 4: A depiction of the Bouncing Trajectory for $\mathcal{C}_{\phi}\text{-}\textsc{MSP}_{1}(1)$. Angle $\theta$ determines the sequence of points $\Gamma_{i}$, as per Lemma 11, using sequence $q_{i}=\left(\tfrac{\sin\left(\theta\right)}{\sin\left(\theta-\phi\right)}\right)^{i}$.
@@
Next, we compute the points $\Gamma_{i}$ in Algorithm 4 and derive the search cost using Lemma 10. The parameter $\beta=\beta(\phi,\theta)$ defines the sequence $\{g_{j}\}_{j}$, which uniquely determines the points $\{\Gamma_{j}\}_{j}$ and vice versa. We now provide a geometric and inductive definition of these points. For reference, Figure 4 illustrates the construction and highlights the role of the angle $\theta$.
@@
Using Lemma 10, the search cost of Algorithm 4 is
@@
By Lemma 11, if $\theta\in(\phi,\pi/2+\phi/2)$, the sequence is valid for the algorithm, and the resulting search cost is